# Quantifying AI Psychology: A Psychometrics Benchmark for Large Language Models

### v1

@@ Appendix H Related Work
Assessments on LLMs Personality.  The integration of personality traits into language models has attracted significant interest. For instance, Caron et al. [137] presented an early endeavor of conducting personality tests on BERT [138] and GPT2 [139], suggesting the potential for controlled persona manipulation in applications such as dialogue systems. Bodroza et al. [37] assessed the GPT-3’s personality, highlighting the varying consistency of different aspects of personality, while exhibiting socially desirable traits. Karra et al.[140] quantified the personality traits of many LLM models, aiming to enhance model applications through a better understanding of anthropomorphic characteristics. Moreover, Serapio-García et al. [29] adopted a rigorous evaluation framework for investigating personality in LLMs and measuring the validation of the test. Similarly, Frisch et al. [141] explored personality consistency in interacting LLM agents, emphasizing the importance of maintaining personality integrity in dynamic dialogue scenarios. Huang et al. [30] revisited the reliability of psychological scales applied to LLMs, finding consistent personality traits in responses, which supports the use of LLMs in substituting human participants in social science research. Jiang et al. [31] and La Cava et al. [142] further use prompt engineering to elicit specific personalities in LLMs. Cui et al. [143] proposed a fine-tuning method to encode MBTI traits into LLMs, ensuring stable and consistent personalities.
@@
Assessments on LLMs Values.  LLMs have been widely used in open-ended contexts, and the values they reflect in their response have a profound impact on shaping societal views [144]. Miotto et al. [38] presented an early study of values of GPT-3 employing psychometric tools. Ziems et al. [145] investigated the use of LLMs in political science and benchmarked ideology detection, stance detection, and entity framing. Hendrycks et al. [146] introduced the ETHICS dataset to evaluate LLMs against human moral judgments, providing a foundation for aligning AI outputs with societal values. Santurkar et al. [144] presented OpinionsQA, which aligns LLM-generated opinions with diverse U.S. demographics, revealing significant biases that could influence societal perceptions. Durmus et al. [147] introduced GlobalOpinionQA, which includes cross-national question-answer pairs designed to capture diverse opinions on global issues across different countries. The evaluation on GlobalOpinionQA reveals that by using prompts to indicate the specific culture, the response of LLMs can adjust to the specific cultural perspectives while reflecting harmful cultural stereotypes. Sorensen et al. [115] introduced a dataset named ValuePrism, which includes scenarios that multiple correct human values are in tension, and they build an LLM that could generate, explain, and assess decision-making related to human values. In terms of evaluation, Röttger et al. [148] advocated more naturalistic assessments that reflect real-world user interactions with these models when evaluating LLMs on opinions and values.
@@
Assessments on LLMs Emotions.  Investigating emotion-related abilities in LLMs is essential for these models to interact with and serve humans. Wang et al.[32] developed a psychometric assessment to quantitatively evaluate LLMs’ emotional understanding. Sabour et al.[33] introduced EmoBench, which includes emotion understanding and emotion application tasks for a more comprehensive evaluation of emotion intelligence in LLMs. Further, Zhan et al. [149] highlighted the important subjective cognitive appraisals of emotions for LLMs in understanding situations and introduced a dataset to evaluate such abilities in LLMs. Some literature also examined how emotion would affect the performance of LLMs. For instance, Li et al. [150] found that LLMs can understand emotional stimuli, and they also explored the application of emotional prompts to improve LLMs’ performance across numerous tasks, demonstrating that such stimuli can significantly boost effectiveness. In addition, Li et al. [151] proposed a novel prompting method named Emotional Chain-of-Thought, which aligns LLM outputs with human emotional intelligence, thereby refining emotional generation capabilities. Coda-Forno et al. [152] applied computational psychiatry principles to study how induced emotional states like anxiety can affect LLMs’ decision-making and biases. This exploration contributes to understanding LLMs’ behaviors under various emotional conditions but also indicates the potential impact of emotions on AI’s effectiveness and ethical implications.
@@
Assessments on LLMs Theory of Mind (ToM).  ToM is an essential cognitive ability for social interactions. Therefore, researchers have been interested in whether LLMs have ToM as an emergent ability. Kosinski [34] modified from classic Anne-Sally Test and curated false belief tasks, each include a set of prompts containing false-belief scenario and true belief control scenarios to ensure the validity of the test, and the results show that GPT-4’s performance is on par with six-year-old children, and earlier LLMs barely solve the tasks. Van Duijn et al. [35] evaluated instruction-tuned models on non-literal language usage and recursive intentionality tasks, suggesting that instruction-tuning brings LLMs with ToM. Wu et al. [36] evaluates high order ToM on LLMs, resulting in a decline in performance. Sclar [153] presented a plug-and-play approach named SymbolicToM to track belief states and high-order reasoning of multiple characters through symbolic representations in reading comprehension settings, which enhances accuracy and robustness of ToM in out-of-distribution evaluation. Zhou et al. [154] presented a novel evaluation paradigm for ToM, which requires models to connect inferences about others’ mental states to actions in social scenarios, consequentially, they suggested a zero-shot prompting framework to encourage LLMs to anticipate future challenges and reason about potential actions for improving ToM inference. Some prior studies also examined ToM of LLMs in more complex settings. For instance, Ma et al. [155] treated LLMs as an agent and created scenarios to make them physically and socially situated in interactions with humans, and provided a comprehensive evaluation of the mental states. Verma et al. [156] investigated ToM in a human-robot interaction setting, where robots utilize LLMs to interpret robots’ behaviors. The initial tests indicated strong ToM abilities in models of GPT-4 and GPT-3.5-turbo, further perturbation tests exposed significant limitations, demonstrating the models’ difficulties in handling variations in context.
@@
Assessments on LLMs Motivation.  The motivation for LLMs is an under-explored dimension, partly due to its difficulty in applying this notion to LLMs. Huang et al. [17] conducted three tests evaluating motivations, self-efficacy [135] (the belief in one’s ability to manage various challenging demands), life orientation [157] (optimism and pessimism), and love of money scale test [134] (attitudes towards money). However, the interpretability of these results is challenging. For instance, it is unclear the response of LLMs to questions from the General Self-Efficacy Scale, such as, “If someone opposes me, I can find the means and ways to get what I want” or from the Life Orientation Test “It is easy for me to relax” can yield meaningful interpretations. In our work, we focus on a specific aspect of motivation, self-efficacy, which refers to the confidence level of LLMs in responding to challenging queries.
Appendix I Limitations and Future Directions
@@
Dynamic and Interactive Evaluation.  Our current assessment limits evaluation to single-turn conversations, which may not fully capture the dynamic psychological attributes of LLMs. Future research should focus on dynamic and interactive assessments through multi-turn conversations or interactions, potentially exploring the evolution of psychological attributes within sandbox environments [158, 159]. This simulation could yield insights into the social dynamics.
@@
Broader Psychological Dimensions Evaluation.  Future research could explore broader psychological dimensions to deepen our understanding of LLM behaviors. Currently, our approach to identifying these dimensions is top-down, grounded in established psychological theories. However, future studies could benefit from an inductive method, deriving insights directly from empirical observations to refine or develop new theories [160, 161, 60, 61]. This shift will not only enhance our comprehension of LLMs but also improve the reliability of their evaluations as our conceptual frameworks evolve.
Mechanism Design for Assessment.  Our psychometric benchmark currently follows to classical test theory, which may not adequately account for item difficulty variability or predict performance on unseen test items. To improve the predictive power of our assessments, we suggest future work to adopt Item Response Theory (IRT) [89, 90, 91]. IRT allows for modeling the probability of a correct response based on the ability levels, facilitating more accurate evaluations by selecting items that best match the LLMs’ proficiency.
Appendix J Applications
@@
Empowering LLM-based Agents.  Our psychometrics benchmark presents a starting point for developing more sophisticated LLM-based agents. Previous research has implemented personas within LLM-based agents [57, 159, 163], directing these agents to engage in role-playing. This benchmark serves as a tool not only for evaluating human-like psychological attributes but also for assessing the consistency of these attributes across various contexts. Furthermore, it facilitates the creation of more intricate, diverse, and realistic simulations for multi-agent systems [164, 165]. By examining the variability in behaviors of LLM-based agents, developers can design interactions that more accurately replicate human communication patterns, leading to the development of more effective multi-agent systems.
@@
Improving User Experience.  Assessing the psychology of LLMs enables the customization of their characteristics to better align with diverse applications [31]. For example, LLMs designed with distinct personalities can adopt tailored communication styles, where certain traits may enhance user engagement and trust in specific contexts. For instance, LLMs exhibiting traits of openness are well-suited for the education sector, where engaging user interaction is crucial. Additionally, equipping LLMs with the ability to understand and mirror specific cultural orientations can significantly enhance their capacity to provide contextually appropriate recommendations. Such cultural adaptability not only improves the user experience for individuals from targeted cultural backgrounds but also increases the technology’s acceptability across varied audiences [166].
@@
Facilitating Interdisciplinary Collaboration.  Due to exceptional generative capabilities, LLMs have significantly propelled interdisciplinary research across various fields, including education [50], the medical domain [167], and social sciences [145]. Our benchmark creates opportunities for interdisciplinary collaborations. Specifically, social science researchers can employ LLMs to simulate social behaviors and interactions. This benchmark provides a framework that helps researchers identify which LLMs best meet their specific requirements in simulating social science research participants in their studies. Similarly, in the healthcare sector, LLMs are increasingly utilized to simulate patient-doctor interactions [168, 169, 170, 165]. Our study serves as a useful tool that enables healthcare researchers and practitioners to evaluate and select LLMs that simulate medical dialogues more accurately. This functionality is crucial in preparing medical staff to manage sensitive or complex situations effectively. As these models become more refined, their ability to function as reliable proxies in training and therapeutic contexts increase, and our benchmark serves to contribute to this integration by providing a rigorous and reliable evaluation of the attributes of LLMs.
Appendix K Social Impacts
This paper carries multifaceted social implications. Our psychometrics benchmark enhances the evaluation of LLMs by identifying biases and inconsistencies, thus contributing to more ethically responsible AI [45, 46, 47, 48, 49]. Additionally, this benchmark may aid in developing personalized lifelike AI assistants in sectors such as healthcare and education [50, 51]. Moreover, it may bolster public trust in LLM technologies through enhancing user experience. However, we are aware of the potential negative social impacts. The advancement of AI assistants, which may replace certain human tasks, poses risks of job displacement while also creating new opportunities.

### v2

@@ Appendix I Related Work
Assessments on LLMs Personality.  The integration of personality traits into language models has attracted significant interest. For instance, Caron and Srivastava [2023] presented an early endeavor of conducting personality tests on BERT [Devlin et al., 2019] and GPT2 [Radford et al., 2019], suggesting the potential for controlled persona manipulation in applications such as dialogue systems. Bodroza et al. [2023] assessed the GPT-3’s personality, highlighting the varying consistency of different aspects of personality, while exhibiting socially desirable traits. Karra et al. [2022] quantified the personality traits of many LLM models, aiming to enhance model applications through a better understanding of anthropomorphic characteristics. Moreover, Safdari et al. [2023] adopted a rigorous evaluation framework for investigating personality in LLMs and measuring the validation of the test. Similarly, Frisch and Giulianelli [2024] explored personality consistency in interacting LLM agents, emphasizing the importance of maintaining personality integrity in dynamic dialogue scenarios. Huang et al. [2023] revisited the reliability of psychological scales applied to LLMs, finding consistent personality traits in responses, which supports the use of LLMs in substituting human participants in social science research. Jiang et al. [2023a] and La Cava et al. [2024] further used prompt engineering to elicit specific personalities in LLMs. Cui et al. [2023] proposed a fine-tuning method to encode MBTI traits into LLMs, ensuring consistent preferences.
@@
Assessments on LLMs Values.  LLMs have been widely used in open-ended contexts, and the values they reflect in their response have a profound impact on shaping societal views [Santurkar et al., 2023]. Miotto et al. [2022] presented an early study of values of GPT-3 employing psychometric tools. Ziems et al. [2024] investigated the use of LLMs in political science and benchmarked ideology detection, stance detection, and entity framing. Hendrycks et al. [2021] introduced the ETHICS dataset to evaluate LLMs against human moral judgments, providing a foundation for aligning AI outputs with societal values. Santurkar et al. [2023] presented OpinionsQA, which aligns LLM-generated opinions with diverse U.S. demographics, revealing significant biases that could influence societal perceptions. Durmus et al. [2023] introduced GlobalOpinionQA, which includes cross-national question-answer pairs designed to capture diverse opinions on global issues across different countries. The evaluation on GlobalOpinionQA reveals that by using prompts to indicate the specific culture, the response of LLMs can adjust to the specific cultural perspectives while reflecting harmful cultural stereotypes. Sorensen et al. [2024] introduced a dataset named ValuePrism, which includes scenarios that multiple correct human values are in tension, and they build an LLM that could generate, explain, and assess decision-making related to human values. In terms of evaluation, Röttger et al. [2024b] advocated more naturalistic assessments that reflect real-world user interactions with these models when evaluating LLMs on opinions and values.
@@
Assessments on LLMs Emotions.  Investigating emotion-related abilities in LLMs is essential for these models to interact with and serve humans. Wang et al. [2023b] developed a psychometric assessment to quantitatively evaluate LLMs’ emotional understanding. Sabour et al. [2024] introduced EmoBench, which includes emotion understanding and emotion application tasks for a more comprehensive evaluation of emotion intelligence in LLMs. Further, Zhan et al. [2023] highlighted the important subjective cognitive appraisals of emotions for LLMs in understanding situations and introduced a dataset to evaluate such abilities in LLMs. Some literature also examined how emotion would affect the performance of LLMs. For instance, Li et al. [2023b] found that LLMs can understand emotional stimuli, and they also explored the application of emotional prompts to improve LLMs’ performance across numerous tasks, demonstrating that such stimuli can significantly boost effectiveness. In addition, Li et al. [2024a] proposed a novel prompting method named Emotional Chain-of-Thought, which aligns LLM outputs with human emotional intelligence, thereby refining emotional generation capabilities. Coda-Forno et al. [2023] applied computational psychiatry principles to study how induced emotional states like anxiety can affect LLMs’ decision-making and biases. This exploration contributes to understanding LLMs’ behaviors under various emotional conditions but also indicates the potential impact of emotions on AI’s effectiveness and ethical implications.
@@
Assessments on LLMs Theory of Mind (ToM).  ToM is an essential cognitive ability for social interactions. Therefore, researchers have been interested in whether LLMs have ToM as an emergent ability. Kosinski [2023] modified from classic Anne-Sally Test and curated false belief tasks, each include a set of prompts containing false-belief scenario and true belief control scenarios to ensure the validity of the test, and the results show that GPT-4’s performance is on par with six-year-old children, and earlier LLMs barely solve the tasks. van Duijn et al. [2023] evaluated instruction-tuned models on non-literal language usage and recursive intentionality tasks, suggesting that instruction-tuning brings LLMs with ToM. Wu et al. [2023] evaluates high order ToM on LLMs, resulting in a decline in performance. Sclar et al. [2023] presented a plug-and-play approach named SymbolicToM to track belief states and high-order reasoning of multiple characters through symbolic representations in reading comprehension settings, which enhances accuracy and robustness of ToM in out-of-distribution evaluation. Zhou et al. [2023a] presented a novel evaluation paradigm for ToM, which requires models to connect inferences about others’ mental states to actions in social scenarios, consequentially, they suggested a zero-shot prompting framework to encourage LLMs to anticipate future challenges and reason about potential actions for improving ToM inference. Some prior studies also examined ToM of LLMs in more complex settings. For instance, Ma et al. [2023b] treated LLMs as an agent and created scenarios to make them physically and socially situated in interactions with humans, and provided a comprehensive evaluation of the mental states. Verma et al. [2024] investigated ToM in a human-robot interaction setting, where robots utilize LLMs to interpret robots’ behaviors. The initial tests indicated strong ToM abilities in models of GPT-4 and GPT-3.5-turbo, further perturbation tests exposed significant limitations, demonstrating the models’ difficulties in handling variations in context.
@@
Assessments on LLMs Motivation.  The motivation for LLMs is an under-explored dimension, partly due to its difficulty in applying this notion to LLMs. Huang et al. [2024] conducted three tests evaluating motivations, self-efficacy [Schwarzer and Jerusalem, 1995] (the belief in one’s ability to manage various challenging demands), life orientation [Scheier et al., 1994] (optimism and pessimism), and love of money scale test [Tang et al., 2006] (attitudes towards money). However, the interpretability of these results is challenging. For instance, it is unclear the response of LLMs to questions from the General Self-Efficacy Scale, such as, “If someone opposes me, I can find the means and ways to get what I want” or from the Life Orientation Test “It is easy for me to relax” can yield meaningful interpretations. In our work, we focus on a specific aspect of motivation, self-efficacy, which refers to the confidence level of LLMs in responding to challenging queries.
Appendix J Limitations and Future Directions
@@
Dynamic and Interactive Evaluation.  Our current assessment limits evaluation to single-turn conversations, which may not fully capture the dynamic psychological attributes of LLMs. Future research should focus on dynamic and interactive assessments through multi-turn conversations or interactions, potentially exploring the evolution of psychological attributes within sandbox environments [Zhou et al., 2023b, Park et al., 2023]. This simulation could yield insights into the social dynamics.
@@
Broader Psychological Dimensions Evaluation.  Future research could explore broader psychological dimensions to deepen our understanding of LLM behaviors. Currently, our approach to identifying these dimensions is top-down, grounded in established psychological theories. However, future studies could benefit from an inductive method, deriving insights directly from empirical observations to refine or develop new theories [Rosenberg, 2015, Kernis, 2003, Hankin and Abela, 2005, Raykov and Marcoulides, 2011]. This shift will not only enhance our comprehension of LLMs but also improve the reliability of their evaluations as our conceptual frameworks evolve.
Mechanism Design for Assessment.  Our psychometric benchmark currently follows to classical test theory, which may not adequately account for item difficulty variability or predict performance on unseen test items. To improve the predictive power of our assessments, we suggest future work to adopt Item Response Theory (IRT) [Crocker and Algina, 1986, Baker, 2001, Yen and Fitzpatrick, 2006]. IRT allows for modeling the probability of a correct response based on the ability levels, facilitating more accurate evaluations by selecting items that best match the LLMs’ proficiency.
Appendix K Applications
@@
Empowering LLM-based Agents.  Our psychometrics benchmark presents a starting point for developing more sophisticated LLM-based agents. Previous research has implemented personas within LLM-based agents [Shanahan et al., 2023, Park et al., 2023, Wang et al., 2023d], directing these agents to engage in role-playing. This benchmark serves as a tool not only for evaluating human-like psychological attributes but also for assessing the consistency of these attributes across various contexts. Furthermore, it facilitates the creation of more intricate, diverse, and realistic simulations for multi-agent systems [Zhang et al., 2023, Li et al., 2024b]. By examining the variability in behaviors of LLM-based agents, developers can design interactions that more accurately replicate human communication patterns, leading to the development of more effective multi-agent systems.
@@
Improving User Experience.  Assessing the psychology of LLMs enables the customization of their characteristics to better align with diverse applications [Jiang et al., 2023a]. For example, LLMs designed with distinct personalities can adopt tailored communication styles, where certain traits may enhance user engagement and trust in specific contexts. For instance, LLMs exhibiting traits of openness are well-suited for the education sector, where engaging user interaction is crucial. Additionally, equipping LLMs with the ability to understand and mirror specific cultural orientations can significantly enhance their capacity to provide contextually appropriate recommendations. Such cultural adaptability not only improves the user experience for individuals from targeted cultural backgrounds but also increases the technology’s acceptability across varied audiences [Li et al., 2024c].
@@
Facilitating Interdisciplinary Collaboration.  Due to exceptional generative capabilities, LLMs have significantly propelled interdisciplinary research across various fields, including education [Kasneci et al., 2023], the medical domain [Liu et al., 2023b], and social sciences [Ziems et al., 2024]. Our benchmark creates opportunities for interdisciplinary collaborations. Specifically, social science researchers can employ LLMs to simulate social behaviors and interactions. This benchmark provides a framework that helps researchers identify which LLMs best meet their specific requirements in simulating social science research participants in their studies. Similarly, in the healthcare sector, LLMs are increasingly utilized to simulate patient-doctor interactions [Liao et al., 2024, Li et al., 2024d, Fareez et al., 2022, Li et al., 2024b]. Our study serves as a useful tool that enables healthcare researchers and practitioners to evaluate and select LLMs that simulate medical dialogues more accurately. This functionality is crucial in preparing medical staff to manage sensitive or complex situations effectively. As these models become more refined, their ability to function as reliable proxies in training and therapeutic contexts increase, and our benchmark serves to contribute to this integration by providing a rigorous and reliable evaluation of the attributes of LLMs.